How would you multiply the SQUARE ROOTS of:
14 x 2

Answer by ptaylor(1894)

To multiply (sqrt14)(sqrt2)
sqrt14=(sqrt(2)(7))
(sqrt2)(sqrt2)=2 so we have
(sqrt(2)(7))(sqrt2)=
(2)(sqrt7)
0r we could do it another way:
(sqrt14)(sqrt2)=sqrt(28)=(sqrt(4)(7))=
(2)sqrt(7)
Hope this helps. Happy holidays----ptaylor